A boat is travelling 4.00 m/s north across a river that is 125 m wide. The river current is moving 2.10 m/s east. What is the actual speed of the boat?


1
Expert's answer
2021-01-10T18:28:23-0500
"V=\sqrt{v^2+v^2_{river}}\\V=\sqrt{4^2+2.1^2}=4.52\frac{m}{s}"